OpenStudy (anonymous):

solve for x. 5^(2x)=17

OpenStudy (helder_edwin):

use logarithms

OpenStudy (anonymous):

yeah how

OpenStudy (helder_edwin):

\[\large 5^{2x}=17 \] \[\large \ln 5^{2x}=\ln(17) \] \[\large 2x\ln(5)=\ln(17) \] \[\large 2x=\frac{\ln(17)}{\ln(5)} \] u can finish this

OpenStudy (anonymous):

i cant:/

OpenStudy (helder_edwin):

really?

OpenStudy (helder_edwin):

well \[\large x=\frac{\ln(17)}{2\ln(5)} \]

OpenStudy (anonymous):

0.4300

OpenStudy (helder_edwin):

i got 0,880187213865

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

5^(6x)=220

OpenStudy (anonymous):

i got 0.2729

OpenStudy (helder_edwin):

it is wrong. do what i did and show me your work

OpenStudy (anonymous):

i did i got x=log 220/ 6log 5

OpenStudy (helder_edwin):

cool

OpenStudy (helder_edwin):

i got 0,558541536424

OpenStudy (anonymous):

how?

OpenStudy (helder_edwin):

first log 220=2,3424 then log 5=0,6989 this times 6 = 4,1938 then 2,3424 divided by 4,1938 = 0,5585

OpenStudy (anonymous):

ohhhh... solve. 7^x=3145

OpenStudy (helder_edwin):

u can do it yourself now, right?